OpenStudy (anonymous):

factor this n^2+n-2250

OpenStudy (anonymous):

Mathematica is unable to factor the expression.

OpenStudy (anonymous):

x=(-1+/- sqrt(9001))/2

OpenStudy (anonymous):

The problem is to factor the expression as I read it. Not to set the expression to zero and solve for n.\[\left\{\left\{n\to \frac{1}{2} \left(-1-\sqrt{9001}\right)\right\},\left\{n\to \frac{1}{2} \left(-1+\sqrt{9001}\right)\right\}\right\} \]

OpenStudy (anonymous):

@ Angel1116 Was your problem request to factor the expression or solve some sort of equation? Thank you in advance.